# Forecasts for Constraining Lorentz-violating Damping of Gravitational Waves from Compact Binary Inspirals

### 13 Feb 2024 v1

@@ I INTRODUCTION
In the theoretical realm, various modified theories of gravity Cognola:2006eg ; Copeland:2006wr ; Frieman:2008sn have been proposed to address challenges within GR. A subset of these theories has garnered significant attention for deviating from a fundamental principle of GR — the Lorentz invariance. At high-energy levels, it is widely believed that this invariance will be broken when gravity is quantized. Various modified gravity theories have been proposed to explore the nature of Lorentz violation in gravity, including the Einstein-Æther theory Jacobson:2000xp ; Eling:2004dk ; Jacobson:2007veq ; Li:2007vz ; Battye:2017zvv ; Zhang:2023kzs ; Liu:2021yev ; Zhu:2019ura , Hořava-Lifshitz theories of quantum gravity Horava:2009uw ; Takahashi:2009wc ; Wang:2012fi ; Zhu:2013fja , the standard model extension Kostelecky:2016kfm ; Bailey:2006fd ; Mewes:2019dhj ; Shao:2020shv ; Wang:2021ctl ; Kostelecky:2017zob ; Niu:2022yhr , and spatial covariant gravities Gao:2020qxy ; Gao:2020yzr ; Gao:2019twq ; Gao:2019liu ; Joshi:2021azw .
The violation of Lorentz symmetry in gravity can introduce deviations from GR in the propagation of GWs. These deviations manifest in two distinct ways, influencing the propagation behavior of GWs in the cosmological background. First, with Lorentz violation, the conventional linear dispersion relation of GWs can be modified into a nonlinear one, which in turn changes the phase velocities of GWs at different frequencies. This effect can arise from a large number of Lorentz-violating theories. Second, Lorentz violation can introduce frequency-dependent friction into the propagation equation of GWs, resulting in varying damping rates for GWs of different frequencies during their propagation. This effect normally arises from those theories with mixed temporal and spatial derivatives of the spacetime metric in the modified theories of gravity with spatial covariance, for example, the Hořava-Lifshitz gravity Colombo:2014lta , the spatial covariant gravities Gao:2019liu ; Zhu:2022uoq , etc.
Testing Lorentz symmetry of gravitational interaction by using the observational data from GW events in LIGO-Virgo-KAGRA catalogs and future GW detectors has been carried out in a lot of works, see Refs. LIGOScientific:2019fpa ; LIGOScientific:2020tif ; Mirshekari:2011yq ; Will:1997bb ; Zhu:2022uoq ; Gong:2021jgg and references therein. In most of these works, the effects due to the nonlinear dispersion relation have been extensively considered. Recently, the constraint on the Lorentz-violating damping effects from GW events in LIGO-Virgo-KAGRA catalogs was first obtained Zhu:2023wci .
@@ II.1 Propagating equation of GWs with Lorentz-violating damping effects
Let us first investigate the propagation of GWs with Lorentz-violating damping effect on a flat Friedmann-Robertson-Walker spacetime. Treating this spacetime as a background, GWs can be described by the tensor perturbations of the metric, where the metric is expressed in the form of
@@
$h_{A}^{\prime\prime}+(2+{\bar{\nu}}+\nu_{A})\mathcal{H}h^{\prime}_{A}+(1+\bar{\mu}+\mu_{A})k^{2}h_{A}=0,$ (2.5)
where a prime denotes the derivative concerning the conformal time $\tau$ and ${\cal H}=a^{\prime}/a$. The four parameters, ${\bar{\nu}}$, $\nu_{A}$, ${\bar{\mu}}$, and $\mu_{A}$ label the new effects on the propagation of GWs arising from theories beyond GR. As mentioned in Ref. Zhu:2023wci , such parametrization provides a general framework for exploring possible modified GW propagations arising from a large number of modified theories of gravity. Different parameters correspond to different effects on the propagation of GWs. The parameters $\nu_{A}$ and $\mu_{A}$ represent the effects of parity violations, while the parameters ${\bar{\nu}}$ and ${\bar{\mu}}$, if frequency-dependent, can originate from other potential modifications involving Lorentz violations. For ${\bar{\nu}}$ and ${\bar{\mu}}$, the former provides an amplitude modulation of the GW waveform, while the latter one ${\bar{\mu}}$ determines the phase velocities of the GWs. In this paper, we will only concentrate on the case of the Lorentz-violating damping effect, and for this case one has
@@
Thus the propagation equation now can be written in the form of
$h^{\prime\prime}_{A}+(2+{\bar{\nu}})\mathcal{H}h^{\prime}_{A}+k^{2}h_{A}=0.$ (2.7)
With the Lorentz-violating damping effect, the only nonzero parameter ${\bar{\nu}}$ is generally frequency dependent. Thus, following Ref. Zhu:2023wci , one can further parametrize ${\bar{\nu}}$ in the form of
@@
where $\beta_{\bar{\nu}}$ is an arbitrary number, $\alpha_{\bar{\nu}}$ is an arbitrary function of time, and $M_{\rm LV}$ denotes the energy scale of the Lorentz violation. The parameters $\alpha_{\bar{\nu}}$ and $\beta_{\bar{\nu}}$ depend on the specific modified theories of gravity.
In this paper, we consider the case with $\beta_{\bar{\nu}}=2$ and derive the corresponding modified waveform of GWs. We then explore the potential constraints on this modified waveform using proposed GW detectors such as aLIGO, CE, ET, Taiji, TianQin, and LISA. It is worth noting that this case can arise from the mixed temporal and spatial derivatives of the spacetime metric in the modified theories of gravity with spatial covariance Gao:2019liu ; Zhu:2022uoq ; Colombo:2014lta , for example, the mixed term in the form of
$\nabla_{k}K_{ij}\nabla^{k}K^{ij}$ (2.9)
@@
can appear in both the spatial covariant gravity Gao:2019liu and Hořava-Lifshitz gravity Colombo:2014lta , where $\nabla_{k}$ denotes the covariant derivative associated with the spatial metric $g_{ij}$ and $K_{ij}$ is the extrinsic curvature tensor. It is also shown that by including mixed derivative terms, the non-protectable Hořava-Lifshitz gravity could be power-counting renormalizable and free of ghosts Colombo:2014lta . One can connect the coupling coefficient $c_{1}^{(2,2)}$ of Eq. (2.9) in modified gravities to the parameters $\alpha_{\bar{\nu}}$ and $M_{\rm LV}$ via
$\frac{\alpha_{\bar{\nu}}(\tau)}{M_{\rm LV}^{2}}=\frac{2c_{1}^{(2,2)}}{M_{\rm Pl}^{2}},$ (2.10)
where $M_{\rm Pl}$ is the reduced Planck mass. Considering the coupling coefficient as a constant, which is equivalent to treating $\alpha_{\bar{\nu}}$ as constant, we can simply set $\alpha_{\bar{\nu}}=1$ by absorbing its order of magnitude into $M_{\rm LV}$.
II.2 Amplitude modulation of GWs with Lorentz-violating damping effect
@@
Here $h_{A}^{\rm GR}$ is the solution of Eq. (2.7) when ${\bar{\nu}=0}$. ${\cal A}_{A}^{\rm GR}$ and $\Phi^{\rm GR}(\tau)$ are the amplitude and phase of $h_{A}^{\rm GR}$ respectively. With this decomposition, $\theta(\tau)$ encodes the correction arising from ${\bar{\nu}}$ which characterizes the Lorentz-violating damping effect.
@@
In GR, the time derivative of the phase $\Phi^{\prime}\sim k$. Here the wavenumber $k$ is connected to the frequency of GWs by $k=2\pi f/a_{0}$. Since the amplitude correction $\theta$ is induced by the expansion of the universe, one has $\theta^{\prime}\sim{\cal H}$ and $\theta^{\prime\prime}\sim{\cal H}^{2}$. Considering $k\gg{\cal H}$ and $\theta^{\prime\prime}\ll\Phi^{\prime}\theta^{\prime}\sim k\theta^{\prime}$, Eq. (II.2) can be simplified into
@@
II.3 Amplitude modification to the waveform of GWs
@@
where ${\cal A}_{\rm GR}$ and $\Psi_{\rm GR}$ represent the amplitude and phase of GWs of a compact binary inspiral signal in GR, and $e^{\delta h_{2}}$ with $\delta h_{2}$ being given by Eq. (II.2) is the amplitude correction to the waveform of GWs in GR. In the post-Newtonian approximation, the amplitude and the phase of GWs in GR can be expressed as Wang:2022yxb
@@ III.2 Noise power spectral density of detector
$P_{\mathrm{acc}}$ $=\left(3\times 10^{-15}\mathrm{~{}m}\mathrm{~{}s}^{-2}\right)^{2}\left(1+\left(\frac{0.4\,\mathrm{mHz}}{f}\right)^{2}\right)$
×$\times\left(1+\left(\frac{f}{8\,\mathrm{mHz}}\right)^{4}\right)\mathrm{Hz}^{-1}.$ (3.9)
@@
$S_{\rm n}(f)$ $=\frac{10}{3L^{2}}\left[S_{\rm x}+\frac{4S_{\rm a}}{(2\pi f)^{4}}\left(1+\frac{10^{-4}\mathrm{~{}Hz}}{f}\right)\right]$
×$\times\left[1+0.6\left(\frac{f}{f_{*}}\right)^{2}\right],$ (3.11)
@@ IV.2 $M_{\rm LV}$ from space-based detectors
For the space-based detectors, including Taiji, TianQin, and LISA, the constraints on $M_{\rm LV}$ are about three orders of magnitude weaker than those from the ground-based detectors. As shown in Table 2, the constraints on $M_{\rm LV}$ is roughly at $\gtrsim 2\times 10^{-24}\ {\rm GeV}$ for single GW event. When considering a joint analysis of 100 simulated GW events, the constraints from all three detectors are roughly at $\gtrsim 5\times 10^{-24}\;{\rm GeV}$. Among the three detectors, Taiji achieves the best result, with $M_{\mathrm{LV}}>5.86\times 10^{-24}$ GeV. The reason why the constraints from the space-based detectors are weaker than those from the ground-based detectors is easy to understand. As one can see from Eq. (II.2), the amplitude correction to the waveform due to the Lorentz-violating damping effect is proportional to the square of the GW frequencies, which implies this effect is more sensitive to the higher GW frequencies, and thus the ground-based detectors can give stronger constraints than space-based detectors. We note that to estimate the number of events within our joint redshift and total mass range, we employ the data and methods from Refs. Wang:2021srv ; Klein:2015hvg .

### 10 Apr 2024 v2

@@ I INTRODUCTION
In the theoretical realm, various modified theories of gravity have been proposed to address challenges within GR, see refs. Cognola:2006eg ; Copeland:2006wr ; Frieman:2008sn ; Jackiw:2003pm ; Yunes:2010yf ; Yagi:2012vf ; Wang:2017brl ; Sefiedgar:2010we ; Crisostomi:2017ugk ; Gao:2014soa ; Gao:2019lpz and references therein. A subset of these theories has garnered significant attention for deviating from a fundamental principle of GR — the Lorentz invariance. At high-energy levels, it is widely believed that this invariance will be broken when gravity is quantized. Various modified gravity theories have been proposed to explore the nature of Lorentz violation in gravity, including the Einstein-Æther theory Jacobson:2000xp ; Eling:2004dk ; Jacobson:2007veq ; Li:2007vz ; Battye:2017zvv ; Zhang:2023kzs ; Liu:2021yev ; Zhu:2019ura , Hořava-Lifshitz theories of quantum gravity Horava:2009uw ; Takahashi:2009wc ; Wang:2012fi ; Zhu:2013fja , and spatial covariant gravities Gao:2020qxy ; Gao:2020yzr ; Gao:2019twq ; Gao:2019liu ; Joshi:2021azw . A phenomenological framework, the standard model extension, has also been extensively studied in the literature for exploring the possible properties of Lorentz violations in the gravitational sector Kostelecky:2016kfm ; Bailey:2006fd ; Mewes:2019dhj ; Shao:2020shv ; Wang:2021ctl ; Kostelecky:2017zob ; Niu:2022yhr .
The violation of Lorentz symmetry in gravity can introduce deviations from GR in the propagation of GWs. These deviations manifest in two distinct ways, influencing the propagation behavior of GWs in the cosmological background. First, with Lorentz violation, the conventional linear dispersion relation of GWs can be modified into a nonlinear one, which in turn changes the phase velocities of GWs at different frequencies. This effect can arise from a large number of Lorentz-violating theories. Second, Lorentz violation can introduce frequency-dependent friction into the propagation equation of GWs, resulting in varying damping rates for GWs of different frequencies during their propagation. This effect normally arises from those theories with mixed temporal and spatial derivatives of the spacetime metric in the modified theories of gravity with spatial covariance, for example, the Hořava-Lifshitz gravity Colombo:2014lta , the spatial covariant gravities Gao:2019liu ; Zhu:2022uoq , etc. Here we would like to note that the possible Lorentz violations could also lead to source-dependence on the speed of GWs Ghosh:2023xes .
Testing Lorentz symmetry of gravitational interaction by using the observational data from GW events in LIGO-Virgo-KAGRA catalogs and future GW detectors has been carried out in a lot of works, see Refs. LIGOScientific:2019fpa ; LIGOScientific:2020tif ; Mirshekari:2011yq ; Will:1997bb ; Zhu:2022uoq ; Gong:2021jgg ; Haegel:2022ymk and references therein. In most of these works, the effects due to the nonlinear dispersion relation have been extensively considered. Recently, the constraint on the Lorentz-violating damping effects from GW events in LIGO-Virgo-KAGRA catalogs was first obtained Zhu:2023wci .
@@ II.1 Propagating equation of GWs with Lorentz-violating damping effects
Let us investigate the propagation of GWs with Lorentz-violating damping effect on a flat Friedmann-Robertson-Walker spacetime. Treating this spacetime as a background, GWs can be described by the tensor perturbations of the metric, where the metric is expressed in the form of
@@
$h_{A}^{\prime\prime}+(2+{\bar{\nu}}+\nu_{A})\mathcal{H}h^{\prime}_{A}+(1+\bar{\mu}+\mu_{A})k^{2}h_{A}=0,$
where a prime denotes the derivative concerning the conformal time $\tau$ and ${\cal H}=a^{\prime}/a$. The four parameters, ${\bar{\nu}}$, $\nu_{A}$, ${\bar{\mu}}$, and $\mu_{A}$ label the new effects on the propagation of GWs arising from theories beyond GR. As mentioned in Ref. Zhu:2023wci , such parametrization provides a general framework for exploring possible modified GW propagations arising from a large number of modified theories of gravity. Different parameters correspond to different effects on the propagation of GWs. The parameters $\nu_{A}$ and $\mu_{A}$ represent the effects of parity violations, while the parameters ${\bar{\nu}}$ and ${\bar{\mu}}$, if frequency-dependent, can originate from other potential modifications involving Lorentz violations. For ${\bar{\nu}}$ and ${\bar{\mu}}$, the former provides an amplitude modulation of the GW waveform, while the latter one ${\bar{\mu}}$ determines the phase velocities of the GWs.
In this paper, we will only concentrate on the case of the Lorentz-violating damping effect, and for this case one has
@@
In general, due to whether the parameter $\bar{\nu}$ is frequency-independent or not, the effects of $\bar{\nu}$ have two possibilities. When $\bar{\nu}$ is frequency-independent and time-dependent, it can be related to a time-dependent Planck mass $M_{*}(t)$ by writing Lagos:2019kds
${\cal H}{\bar{\nu}}=H\frac{d\ln M_{*}^{2}}{\ln a}.$ (2.7)
See ref. Zhu:2022uoq as well for a specific example with an explicit action for nonzero $\bar{\nu}$ and its relation to the running of the Planck mass $M_{*}(t)$. Another possibility corresponds to a frequency-dependent $\bar{\nu}$, which represents the Lorentz-violating damping effect we studied in this paper. For this case, following Ref. Zhu:2023wci , one can further parametrize ${\bar{\nu}}$ in the form of
@@
where $\beta_{\bar{\nu}}$ is an arbitrary number, $\alpha_{\bar{\nu}}$ is an arbitrary function of time, and $M_{\rm LV}$ denotes the energy scale of the Lorentz violation  [^1] In ref. Zhao:2019xmm , a different symbol $M_{\rm PV}$ is used to represent the Lorentz-violating energy scale. Note that $M_{\rm LV}^{-2}$ in the above parametrization is also directly related to the coefficient ${\cal G}_{2}/{\cal G}_{0}$ used for parametrizing the modified GW propagarions in ref. Gao:2019liu . . The parameters $\alpha_{\bar{\nu}}$ and $\beta_{\bar{\nu}}$ depend on the specific modified theories of gravity. This case can arise from the mixed temporal and spatial derivatives of the spacetime metric in the modified theories of gravity with spatial covariance Gao:2019liu ; Zhu:2022uoq ; Colombo:2014lta . In the next subsection, we present a specific example that induces the Lorentz-violating damping effect in the propagation of GWs.
II.2 A specific example with Lorentz-violating damping effect
@@
To illustrate the Lorentz-violating damping effects clearly, let us consider a specific example, with the mixed term $\nabla_{k}K_{ij}\nabla^{k}K^{ij}$, which can appear in both the spatial covariant gravity Gao:2019liu and Hořava-Lifshitz gravity Colombo:2014lta , where $\nabla_{k}$ denotes the covariant derivative associated with the spatial metric $g_{ij}$ and $K_{ij}$ is the extrinsic curvature tensor. It is also shown that by including mixed derivative terms, the non-protectable Hořava-Lifshitz gravity could be power-counting renormalizable and free of ghosts Colombo:2014lta . With this mixed term, one can write down the general action of the gravitational part with spatial covariance in the form of Zhu:2022uoq
[…]
where the first term represents the Einstein-Hilbert action of GR in the $3+1$ form, the second term signifies one of the modifications to GR, $c_{1}$ is the coupling coefficient which is a function of the lapse function $N$ and time $t$, and $M_{\rm Pl}$ is the reduced Planck mass. Here we would like to mention that in the second term of the action, we also include $R_{ij}R^{ij}$ to eliminate the effect of the mixed derivative term $\nabla_{k}K_{ij}\nabla^{k}K^{ij}$ in the disperation of GWs, such that the GWs propagate at the speed of light. Note that in writting the above action, we adopt the Arnowitt-Deser-Misner (ADM) form ADM . In eq.(LABEL:action_ADM), $N$ is the lapse function, $\nabla_{k}$ denotes the covariant derivative associated with the spatial metric $g_{ij}$ and $K_{ij}$ is the extrinsic curvature tensor in the ADM form. Then, the action of GWs with the $c_{1}$ term up to the quadratic order can be written in the form Gao:2019liu ,
$S^{(2)}$ $=$ $\frac{M_{\rm Pl}^{2}}{8}\int dtd^{3}xa^{3}\left(\dot{h}_{ij}\dot{h}^{ij}+{h}_{ij}\frac{\triangle}{a^{2}}{h}^{ij}\right.$
$\left.-c_{1}\dot{h}_{ij}\frac{\triangle}{a^{2}}\dot{h}^{ij}+c_{1}h_{ij}\frac{\triangle^{2}}{a^{2}}h^{ij}\right).$
Here $\triangle\equiv\delta^{ij}\partial_{i}\partial_{j}$ with $\delta^{ij}$ being the Kronecker delta and a dot denotes the derivative respect to the cosmic time $t$.
Variation of the quadratic action (II.2) with respect to $h_{ij}$, one obtains the equation of motion for $h_{ij}$ as
$\left(1-c_{1}\frac{\partial^{2}}{a^{2}}\right)h^{\prime\prime}_{ij}+\left[2\mathcal{H}-c_{1}^{\prime}\frac{\partial^{2}}{a^{2}}\right]h^{\prime}_{ij}$
$-\left(1-c_{1}\frac{\partial^{2}}{a^{2}}\right)\partial^{2}h_{ij}=0.$ (2.11)
Then using the Fourier harmonics (2.3), the above equation can be cast into the form of eq. (II.1) as
$h^{\prime\prime}_{A}+(2+{\bar{\nu}})\mathcal{H}h^{\prime}_{A}+k^{2}h_{A}=0,$ (2.12)
where
${\cal H}\bar{\nu}=\left[\ln\left(1+c_{1}\frac{k^{2}}{a^{2}}\right)\right]^{\prime}.$ (2.13)
Considering that the effect from $c_{1}$ term is small, one approximately has
${\cal H}\bar{\nu}\simeq\left(c_{1}\frac{k^{2}}{a^{2}}\right)^{\prime}.$ (2.14)
Then, one can connect the coupling coefficient $c_{1}$ in the action (LABEL:action_ADM) to the parameters $\alpha_{\bar{\nu}}$ and $M_{\rm LV}$ via
$c_{1}(\tau)=\frac{\alpha_{\bar{\nu}}(\tau)}{M_{\rm LV}^{2}},$ (2.15)
with $\beta_{\bar{\nu}}=2$.
In this paper, we consider the case with $\beta_{\bar{\nu}}=2$ and derive the corresponding modified waveform of GWs. We then explore the potential constraints on this modified waveform using proposed GW detectors such as aLIGO, CE, ET, Taiji, TianQin, and LISA. This case is induced by $\nabla_{k}K_{ij}\nabla^{k}K^{ij}$ which contains two time derivatives and two spatial derivatives. The cases with $\beta_{\bar{\nu}}>2$ are also possible if one added terms with two time derivatives and more than two spatial derivatives in the gravitational action. However, these higher spatial derivative terms are expected to be suppressed, comparing to the leading-order case with $\beta_{\bar{\nu}}=2$. Therefore, in this paper, we only focus on the leading order one with $\beta_{\bar{\nu}}=2$.
II.3 Amplitude modulation of GWs with Lorentz-violating damping effect
@@
Here $h_{A}^{\rm GR}$ is the solution of Eq. (2.12) when ${\bar{\nu}=0}$. ${\cal A}_{A}^{\rm GR}$ and $\Phi^{\rm GR}(\tau)$ are the amplitude and phase of $h_{A}^{\rm GR}$ respectively. With this decomposition, $\theta(\tau)$ encodes the correction arising from ${\bar{\nu}}$ which characterizes the Lorentz-violating damping effect. We would like to mention that, to obtain a waveform model with the propagation effects due to both the Lorentz-violating damping effect, we assume that the waveform extracted in the binary’s local wave zone is well-described by a waveform in GR. The same assumption has also been used in the analysis for testing the propagation effects in LIGOScientific:2019fpa ; LIGOScientific:2020tif . In this way, one can calculate both the amplitude and phase corrections due to the propagation effects to the GR-based waveform by using the stationary phase approximation (SPA) during the inspiral phase of the binary system Zhao:2019xmm .
@@
In GR, the time derivative of the phase $\Phi^{\prime}\sim k$. Here the wavenumber $k$ is connected to the frequency of GWs by $k=2\pi f/a_{0}$. Since the amplitude correction $\theta$ is induced by the expansion of the universe, one has $\theta^{\prime}\sim{\cal H}$ and $\theta^{\prime\prime}\sim{\cal H}^{2}$. Considering $k\gg{\cal H}$ and $\theta^{\prime\prime}\ll\Phi^{\prime}\theta^{\prime}\sim k\theta^{\prime}$, Eq. (II.3) can be simplified into
@@
II.4 Amplitude modification to the waveform of GWs
@@
where ${\cal A}_{\rm GR}$ and $\Psi_{\rm GR}$ represent the amplitude and phase of GWs of a compact binary inspiral signal in GR, and $e^{\delta h_{2}}$ with $\delta h_{2}$ being given by Eq. (II.3) is the amplitude correction to the waveform of GWs in GR. In the post-Newtonian approximation, the amplitude and the phase of GWs in GR can be expressed as Wang:2022yxb
@@ III.2 Noise power spectral density of detector
$P_{\mathrm{acc}}$ $=$ $\left(3\times 10^{-15}\mathrm{~{}m}\mathrm{~{}s}^{-2}\right)^{2}\left(1+\left(\frac{0.4\,\mathrm{mHz}}{f}\right)^{2}\right)$ (3.9)
×$\times\left(1+\left(\frac{f}{8\,\mathrm{mHz}}\right)^{4}\right)\mathrm{Hz}^{-1}.$
@@
$S_{\rm n}(f)$ $=$ $\frac{10}{3L^{2}}\left[S_{\rm x}+\frac{4S_{\rm a}}{(2\pi f)^{4}}\left(1+\frac{10^{-4}\mathrm{~{}Hz}}{f}\right)\right]$ (3.11)
×$\times\left[1+0.6\left(\frac{f}{f_{*}}\right)^{2}\right],$
@@ IV.2 $M_{\rm LV}$ from space-based detectors
For the space-based detectors, including Taiji, TianQin, and LISA, the constraints on $M_{\rm LV}$ are about three orders of magnitude weaker than those from the ground-based detectors. As shown in Table 2, the constraints on $M_{\rm LV}$ is roughly at $\gtrsim 2\times 10^{-24}\ {\rm GeV}$ for single GW event. When considering a joint analysis of 100 simulated GW events, the constraints from all three detectors are roughly at $\gtrsim 5\times 10^{-24}\;{\rm GeV}$. Among the three detectors, Taiji achieves the best result, with $M_{\mathrm{LV}}>5.86\times 10^{-24}$ GeV. The reason why the constraints from the space-based detectors are weaker than those from the ground-based detectors is easy to understand. As one can see from Eq. (II.3), the amplitude correction to the waveform due to the Lorentz-violating damping effect is proportional to the square of the GW frequencies, which implies this effect is more sensitive to the higher GW frequencies, and thus the ground-based detectors can give stronger constraints than space-based detectors. We note that to estimate the number of events within our joint redshift and total mass range, we employ the data and methods from Refs. Wang:2021srv ; Klein:2015hvg .